OpenStudy (anonymous):

All Calculus students dream :) Find all pair of non constant functions f, g such that (fg)'=f'g' and (fg)''=f''g''

OpenStudy (anonymous):

i saw this yesterday:)

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

watchmath had answered:)

OpenStudy (anonymous):

no he didnt give A,B,C

OpenStudy (anonymous):

i forgot his solution, but

OpenStudy (anonymous):

lets do it now then..:)

OpenStudy (anonymous):

we had 2 eqns. (f/f')+(g/g')=1 and (f'/f")+(g'/g")=1

OpenStudy (anonymous):

its impossible

OpenStudy (anonymous):

we have f'g + fg' = f'g'

OpenStudy (anonymous):

and f''g + 2f'g' + fg'' = f''g''

OpenStudy (anonymous):

ya so dvide by f'g' both sides...tats same!

OpenStudy (nowhereman):

It's easy, just choose functions with disjoint support. Then either f or g and thus f' or g' likewise are zero.

OpenStudy (anonymous):

whats disjoint support

OpenStudy (nowhereman):

Support is the set where they are non-zero. So that just means at every point either of the functions is zero.

OpenStudy (anonymous):

f'g + fg' = f'g', f' ( g' - g ) /g' = f

OpenStudy (anonymous):

we have to find nonconstant solutions nowhereman!!

OpenStudy (anonymous):

f = fg' / ( g' - g)

OpenStudy (nowhereman):

Does f'g' have to be non-constant too?

OpenStudy (anonymous):

my bad f' = fg' / ( g' - g)

OpenStudy (anonymous):

now, take the derivative of this

OpenStudy (anonymous):

cantorset wat r u tryin t say? its f/f'+g/g'=1

OpenStudy (anonymous):

now do you get that?

OpenStudy (anonymous):

what is the problem the 2 eqns. (f/f')+(g/g')=1 and (f'/f")+(g'/g")=1

OpenStudy (anonymous):

saub, we know that (fg)' = f'g + fg' , correct ? so f'g + fg' = f'g'

OpenStudy (anonymous):

yes so why are my equations impossible? (f/f')+(g/g')=1 and (f'/f")+(g'/g")=1

OpenStudy (anonymous):

(g/g')+(f/f')=1 fg'' + 2g'f' + f''g=f''g'' saubhik: how do you get this? (f'/f")+(g'/g")=1

OpenStudy (nowhereman):

Well I would substitute \(F = \ln f\) and \(G = \ln g\) so that you get \(1 = F'(1 - 1/G')\) and thus \(G' = \int \frac{1}{1-\frac{1}{G'}}\)

OpenStudy (anonymous):

k...amogh u got me... (f/f')+(g/g')=1 is correct and u have to differentiate this and use the 2nd result...

OpenStudy (anonymous):

Yes and I get this: fg'' + 2g'f' + f''g = f''g''

OpenStudy (anonymous):

ok..so we have to solve: (g/g')+(f/f')=1 and fg'' + 2g'f' + f''g = f''g''

OpenStudy (anonymous):

PS:If this question were my dream it wud be my nightmare :P lol

OpenStudy (anonymous):

Yea but how did you get this? (f'/f")+(g'/g")=1, sorry if its stupid.

OpenStudy (anonymous):

hmmm, ln g(x) + ln f(x) = integral 1 ?

OpenStudy (anonymous):

f/f'+g/g'=1, if you take integral of both sides

OpenStudy (anonymous):

no tats not right i think!

OpenStudy (anonymous):

oh, thats f'/f , woops

OpenStudy (anonymous):

integral of f/f' is not ln f!

OpenStudy (anonymous):

yes..amogh is correct

OpenStudy (anonymous):

Solving these will require deep math cantorset!!

OpenStudy (anonymous):

unless and until u start guessing:P

OpenStudy (anonymous):

it will be ugly to start solving there will be innumerable veriables and integrals...we have to use the exactness condition..

OpenStudy (anonymous):

f=e^{3x/2},g=e^{3x} works . (source:watchmath)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

i remember that solution

OpenStudy (anonymous):

so can we generalize...

OpenStudy (anonymous):

Can you send me the link to the question you posted earlier yesterday?

OpenStudy (anonymous):

Thanks

OpenStudy (nowhereman):

I already told you how to do it. Use the substitution G = ln g and F = ln f or in other terms g = e^G and f = e^F, then for every G you can find an F so that f'g' = (fg)' because the condition can be reformulated to say F' = 1/(1-1/G') which can be solved by integration.

OpenStudy (anonymous):

youre missing a G there

OpenStudy (anonymous):

what is integral(G'/(G'-1)) nowhereman?

OpenStudy (anonymous):

nowhereman is not missing a G cantorset..he used 1/(1-1/G') but how to do this integral???

OpenStudy (anonymous):

, his G is not the same as g

OpenStudy (anonymous):

G = ln g and F = ln f

OpenStudy (anonymous):

yups..i know.....but he is right cantorset u r a calculus tutor..how do we do this integral? any ideas?

OpenStudy (anonymous):

F'=G'/(G'-1) is correct.

OpenStudy (anonymous):

it is integral G ' / ( 1- G' ) ?

OpenStudy (anonymous):

F'=G'/(G'-1) i mean

OpenStudy (anonymous):

yups..integrate both sides u get F=integral(G'/(G'-1)) what will be the RHS?

OpenStudy (anonymous):

first divide that

OpenStudy (anonymous):

k. so 1/(1-1/G') then?

OpenStudy (anonymous):

G'/(G'-1) = 1 + 1/ G' -1

OpenStudy (anonymous):

G'/(G'-1) = 1 + 1/ (G' -1)

OpenStudy (anonymous):

how will u do the later? its the same problem :P

OpenStudy (anonymous):

hmmm, with respect to x ?

OpenStudy (anonymous):

its a differential equation...we can differentiate with respect to the variables of the function..i.e. G(x) wrt x and F(x) wrt x...

OpenStudy (anonymous):

ok...lets go to the main problem after making nowhereman's substitutions we get F'G'=F'+G'..how will u solve this diff eqn?

OpenStudy (anonymous):

dunno

OpenStudy (anonymous):

please dont make fun of me...see that F'G'=F'+G' implies integral(dF(x)*dG(x))=F(x)+G(x) thats an integration with respect to 2 variables!!!

OpenStudy (anonymous):

we need multiple integration here? :)

OpenStudy (anonymous):

am here cantorset..

OpenStudy (anonymous):

ok , im lost, how did you get that , capital F' and G'

OpenStudy (anonymous):

see F=lnf and G=lng then use f'g'=(fg)'

OpenStudy (anonymous):

itll give F'G'=F'+G'

OpenStudy (anonymous):

then integrate but i dont know the integral of LHS..

OpenStudy (watchmath):

f'g+fg'=f'g' (f'-f)g'-f'g=0 g'- f'g/(f'-f)=0 Now you can use integrating factor method. The answer will be still in terms of f and f'. But that is ok, we will use the second requirement.

OpenStudy (anonymous):

yah...

OpenStudy (anonymous):

ok , multiply both sides by e ^ int ( f' / (f' - f)

OpenStudy (watchmath):

yes :)

OpenStudy (anonymous):

what is int ( f' / (f' - f)

OpenStudy (anonymous):

keep it like this? we need to simplify...

OpenStudy (watchmath):

just keep it like that

OpenStudy (anonymous):

g.e ^ int ( f' / (f' - f) =0 is the answer

OpenStudy (watchmath):

=constant

OpenStudy (anonymous):

yess..

OpenStudy (anonymous):

how will u use the 2nd fact watchmath?

OpenStudy (anonymous):

this is pretty ugly to differentiate..

OpenStudy (watchmath):

just plugin the g. You need fundamental calculus theorem to compute g'.

OpenStudy (anonymous):

website keeps frezing on me

OpenStudy (anonymous):

ok im a little slow, one sec

OpenStudy (anonymous):

watchmath how will u extract the g?

OpenStudy (anonymous):

g'- f'/(f'-f)*g =0

OpenStudy (anonymous):

k...lot of work though..

OpenStudy (anonymous):

but process is clear..

OpenStudy (anonymous):

now we know , dy/dx + p(x) *y = q(x), the solution is y = [int ( int e^p(x) )*q(x) dx + c ] / e^int p(x)

OpenStudy (anonymous):

g.e^[f'/(f'-f)]=constant.

OpenStudy (anonymous):

now since q(x) is zero, that simplifies the numerator, so its g = C / ( e^ int ( f

OpenStudy (anonymous):

ln|g|+[f'/(f'-f)]=constant